The supplement of an angle is 30 more than twice the complement of the same angle. What's the angle's measure?
Answer by oberobic(2304)

a = angle
s = supplement of angle a
s = 180-a
c = complement of angle a
c = 90-a
.
s = 2c + 30
.
s = s, so
.
180-a = 2c + 30
.
180-a = 2(90-a) + 30
.
180 -a = 180 -2a + 30
.
a = 30
s = 180-30 = 150
c = 90-30 = 60
.
Check the relationship of s and c.
s = 2c + 30 ??
2c + 30 = 2(60)+30 = 120+30 = 150
correct
.
Angle = 30 degrees.
.
Done.